You recently received a job with a company that manufactures an automobile antitheft device. To conduct an advertising campaign for the product, you need to make a claim about the number of automobile thefts per year. Since the population of various cities in South Africa varies, you decide to use rates per 10,000 people. (The rates are based on the number of people living in the cities.) Your boss said that last year the theft rate per 10,000 people was 44 vehicles. You want to see if it has changed. The following are rates per 10,000 people for 36 randomly selected locations in the South Africa.

55 42 125 62 134 73 39 69 23 94 73 24 51 55 26 66 41 67 15 53 56 91 20 78 7025 62 115 17 36 58 56 33 75 20 16

Using this information, answer the following questions. 1. What critical value(s) would you use?



1
Expert's answer
2021-11-16T15:32:44-0500

we use critical value for two-tailed t test:

for "\alpha=0.05" and "df=n-1=36-1=35" :

"t_{crit}=2.030"